# Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods

### v3

@@ 3.6 Model 6
Our sixth data simulation model is
@@
where $P=0.16$ and $M_{0}=0$. The coefficients $c_{1}=0.3655$, $c_{2}=0.7310$ and $c_{3}=0.3000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$ and $t_{\mathrm{1,max,2}}$ values given in Table 6 (Column 1). This simulated data sample is not “too short” because $P_{1}<\Delta T$ (Equation 27). There is no trend because $p(t)=M_{0}=0$ (Equation 28). These simulated data contain only one signal (Equation 29). However, this simulation Model 6 is not a $K_{2}=1\texteuro$ pure sine model (Equation 30). This double wave simulation model has two unequal minima and maxima. Its DCM orders are $K_{1}=1$, $K_{2}=2$ and $K_{3}=0$. We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
The DCM time series analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 6 (Columns 2-4). The DCM detects the correct $P_{1}$, $A_{1}$, $t_{\mathrm{1,min,1}}$ $t_{\mathrm{1,min,2}}$ $t_{\mathrm{1,max,1}}$ $t_{\mathrm{1,max,2}}$ and $M_{0}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination. We demonstrate the DCM analysis results for simulated data having $n=50$ and ${\mathrm{SN}}=50$ (Figures 6a-d). The long and short searches give $P_{1}=0.159$ (Figure 6a: diamond) and $P_{1}=0.160$ (Figure 6b: diamond). The DCM model $g(t)$ black line covers the black dots denoting the data $y_{i}$ (Figure 6c). The "trend" at $p(t)=M_{0}=-0.0055\pm 0.0026$ is correct. The detrended model g(t)p(t)g(t)-p(t) (black continuous line), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the detrended data $y(t_{i})-p(t_{i})$ (black dots) are displayed in Figure 6d. The thick continuous red line stays under the thin continuous black line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -1.8 (blue dotted line). These residuals show no trends and their level is stable.
@@
We conclude that the DFT “detects” the $P_{1}/2$ and $P_{1}$ periods, where $P_{1}$ is the correct simulated period value. However, the DFT two signal model is not the correct model for these Model 6 simulated data, which contain only one signal. If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$.
The DCM analysis succeeds for Model 6 simulated data. The DFT analysis fails.
@@ 3.7 Model 7
In this section, we analyse our most complex data set. The seventh simulation model is
@@
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$, $P_{1}=1.2$, $P_{2}=1.4$, $M_{0}=1$ $M_{1}=0.25$ and $M_{2}=0.5$. The coefficients $c_{1}=0.3687$, $c_{2}=0.7374$, $c_{3}=0.4000$, $c_{4}=0.3708$, $c_{5}=0.7416$ and $c_{6}=0.6000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{1,max,2}}$, $A_{2}$, $t_{\mathrm{2,min,1}}$, $t_{\mathrm{2,min,2}}$, $t_{\mathrm{2,max,1}}$ and $t_{\mathrm{2,max,2}}$ values given in Table 7 (Column 1). This simulated data sample is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 27). The parabolic $p(t)$ represents “a trend” (Equation 28). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” because $\Delta f=f_{1}-f_{2}=0.12<f_{0}=1$ (Equation 29). The two $h_{1}(t)$ and $h_{2}(t)$ signals are not “pure sines” (Equation 30). These signals are double waves having two unequal minima and maxima. All four main reasons that can cause the failure of DFT analysis are present (Equations 27-30). Therefore, this simulated data set is the most complex one of all analysed seven data sets. Our DCM and DFT time series analysis of Model 7 simulated data is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
@@
The DCM orders of Model 7 are $K_{1}=2$, $K_{2}=2$ and $K_{3}=2$. This model has $\eta=13$ free parameters (Equation 8). We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 7 (Columns 2-4). The DCM analysis results are displayed for one sample of Model 7 simulated data (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$). Since there are $\eta=13$ free model parameters, this $n=100$ sample is quite small for time series analysis. The long and short DCM searches give $P_{1}=1.200$ and $P_{2}=1.385$ (Figure 7a: diamonds), and $P_{1}=1.199$ and $P_{2}=1.405$ (Figure 7b: Diamonds). The black $g(t)$ model line goes through all black $y_{i}$ data dots (Figure 7c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.950\pm 0.050$, $M_{1}=0.2447\pm 0.0067$ and $M_{2}=0.551\pm 0.050$ (Figure 7c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) are displayed in Figure 7d. The $n=100$ residuals (blue dots) are offset to the level of -1.8 (dotted blue line). These DCM model residuals are small and their level is stable. These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small data sets $(n=100)$, if these data are extremely accurate $({\mathrm{SN}}=5~000~000).$
@@
The DFT time series analysis gives the wrong periods for the original data (Figure 7e: diamond at $P_{1}=0.542$ and for the residuals (Figure 7f: diamond at $P_{2}=1.059$. The DFT also gives wrong $p(t)$ trend coefficients $M_{0}=0.670$, $M_{1}=-0.025$ and $M_{2}=0.406$ (Figure 7g: dashed black line). The DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line) deviates from the data $y_{i}$ (black dots). This deviation is largest at the beginning and the end of the simulated data sample. In our Figure 7h, the black dots denote the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$, the continuous black line denotes the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ and the continuous thick red line denotes pure sine signal $s_{\mathrm{y,DFT}}(t)$ detected from the original data. The DFT model $g_{\mathrm{DFT}}(t)$ gives very low amplitude for the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ detected from the residuals (continuous thinner blue line). The correct simulated peak to peak amplitude for this second signal is much larger, $A_{2}=2.$ The DFT model residuals (blue dots) are offset to -1.5 level and show strong trends.
@@
The DCM time series analysis succeeds for Model 7 simulated data. The DFT time series analysis fails, as predicted by Equations 27-30.
@@
Table 8: Fisher tests. DCM analysis of Model 1 simulated data ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
@@ 3.8 Best model
The $K_{1}$, $K_{2}$ and $K_{3}$ orders of the best model are not necessarily known when some time series analysis method is applied to the real data. We know a priori the best DCM and DFT model orders for the simulated data of Models 1-7 (Sections 3.1-3.7). It could be argued that our DCM analysis succeeds only for this reason.
@@
Larger $F_{R}$ or $F_{\chi}$ values have smaller $Q_{F}$ critical levels. Hence, the probability for the $H_{0}$ hypothesis rejection increases when $F_{R}$ or $F_{\chi}$ increases. If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease. This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ increase (Equations 38 and 39: first terms). At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 38 and 39: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values. This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
@@
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated data (electronic data file Model1n100SN100). (1) Polynomial model. (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8. 33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.

### v4

@@ 3.6 Model 6
Our sixth time series simulation model is
@@
where $P=0.16$ and $M_{0}=0$. The coefficients $c_{1}=0.3655$, $c_{2}=0.7310$ and $c_{3}=0.3000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$ and $t_{\mathrm{1,max,2}}$ values given in Table 6 (Column 1). This simulated time series is not “too short” because $P_{1}<\Delta T$ (Equation 35). There is no trend because $p(t)=M_{0}=0$ (Equation 36). This simulated time series contains only one signal (Equation 37). However, this simulation Model 6 is not a $K_{2}=1\texteuro$ pure sine model (Equation 38). This double wave simulation model has two unequal minima and maxima. Its DCM orders are $K_{1}=1$, $K_{2}=2$ and $K_{3}=0$. We perform the DCM and DFT time series analysis between $P_{\mathrm{min}}=P_{1}/3=0.053$ and $P_{\mathrm{max}}=3P_{1}=0.480$.
The DCM time series analysis results for different $n$ and $\mathrm{SN}$} combinations are given in Table 6 (Columns 2-4). The DCM detects the correct $P_{1}$, $A_{1}$, $t_{\mathrm{1,min,1}}$ $t_{\mathrm{1,min,2}}$ $t_{\mathrm{1,max,1}}$ $t_{\mathrm{1,max,2}}$ and $M_{0}$ values even for the lowest $n=50$ and ${\mathrm{SN}}=10$ combination. For increasing $n$ and $\mathrm{SN}$}, the results for the model parameters converge to correct values (WD-effect).
We demonstrate the DCM analysis results for simulated time series having $n=50$ and ${\mathrm{SN}}=50$ (Figures 6a-d). The long and short searches give $P_{1}=0.159$ (Figure 6a: diamond) and $P_{1}=0.160$ (Figure 6b: diamond). The DCM model $g(t)$ black line covers the black dots denoting the data $y_{i}$ (Figure 6c). The "trend" at $p(t)=M_{0}=-0.0055\pm 0.0026$ is correct. The detrended model g(t)p(t)g(t)-p(t) (black continuous line), the pure sine signal $h_{1}(t)$ (red thick continuous line) and the detrended data $y(t_{i})-p(t_{i})$ (black dots) are displayed in Figure 6d. The thick continuous red line stays under the thin continuous black line because $h_{1}(t)=g(t)-p(t)$. The DCM residuals (blue dots) are offset to the level of -1.8 (blue dotted line). These residuals show no trends and their level is stable.
@@
We conclude that the DFT “detects” the $P_{1}/2$ and $P_{1}$ periods, where $P_{1}$ is the correct simulated period value. However, the DFT two signal model is not the correct model for these Model 6 simulated time series, which contains only one signal. If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$.
The DCM analysis succeeds for Model 6 simulated time series. The DFT analysis fails.
@@ 3.7 Model 7
In this section, we analyse our most complex time series. The seventh simulation model is
@@
where $T=[2(t-t_{\mathrm{mid}})]/\Delta T$, $P_{1}=1.2$, $P_{2}=1.4$, $M_{0}=1$ $M_{1}=0.25$ and $M_{2}=0.5$. The coefficients $c_{1}=0.3687$, $c_{2}=0.7374$, $c_{3}=0.4000$, $c_{4}=0.3708$, $c_{5}=0.7416$ and $c_{6}=0.6000$ determine the $A_{1}$, $t_{\mathrm{1,min,1}}$, $t_{\mathrm{1,min,2}}$, $t_{\mathrm{1,max,1}}$, $t_{\mathrm{1,max,2}}$, $A_{2}$, $t_{\mathrm{2,min,1}}$, $t_{\mathrm{2,min,2}}$, $t_{\mathrm{2,max,1}}$ and $t_{\mathrm{2,max,2}}$ values given in Table 7 (Column 1). This simulated time series is “too short” because $\Delta T<P_{1}<P_{2}$ (Equation 35). The parabolic $p(t)$ represents “a trend” (Equation 36). The signal frequencies $f_{1}$ and $f_{2}$ are “too close” because $\Delta f=f_{1}-f_{2}=0.12<f_{0}=1$ (Equation 37). The two $h_{1}(t)$ and $h_{2}(t)$ signals are not “pure sines” (Equation 38). These signals are double waves having two unequal minima and maxima. All four main reasons that can cause the failure of DFT analysis are present (Equations 35-38). Therefore, this simulated time series is the most complex one of all analysed seven time series. Our DCM and DFT time series analysis of Model 7 simulated time series is performed between $P_{\mathrm{min}}=P_{1}/3=0.4$ and $P_{\mathrm{max}}=3P_{1}=3.6$.
@@
The DCM orders of Model 7 are $K_{1}=2$, $K_{2}=2$ and $K_{3}=2$. This model has $\eta=13$ free parameters (Equation 8). We give the DCM analysis results for different $n$ and $\mathrm{SN}$} combinations in Table 7 (Columns 2-4). The DCM analysis results are displayed for one sample of Model 7 simulated time series (Figures 7a-d: $n=100$ and ${\mathrm{SN}}=5~000~000$). Since there are $\eta=13$ free model parameters, this $n=100$ time series is quite small for time series analysis. The long and short DCM searches give $P_{1}=1.200$ and $P_{2}=1.385$ (Figure 7a: diamonds), and $P_{1}=1.199$ and $P_{2}=1.405$ (Figure 7b: Diamonds). The black $g(t)$ model line goes through all black $y_{i}$ data dots (Figure 7c). The DCM detects the correct polynomial trend $p(t)$ coefficients $M_{0}=0.950\pm 0.050$, $M_{1}=0.2447\pm 0.0067$ and $M_{2}=0.551\pm 0.050$ (Figure 7c: dashed black line). The detrended DCM model g(t)p(t)g(t)-p(t) (white continuous line), the detrended data $y(t_{i})-p(t_{i})$ (black dots), the signal $h_{1}(t)$ (red thick continuous line) and the signal $h_{2}(t)$ (blue thin continuous line) are displayed in Figure 7d. The $n=100$ residuals (blue dots) are offset to the level of -1.8 (dotted blue line). These DCM model residuals are small and their level is stable. These results confirm that the DCM time series analysis method can detect complex non-linear models $(\eta=13)$ from very small time series $(n=100)$, if the data are extremely accurate $({\mathrm{SN}}=5~000~000).$ These results demonstrate the power of WD-effect because both periods $P_{1}$ and $P_{2}$ are shorter than the sample window $\Delta T$, the trend is a parabola and the signals are not pure sines.
@@
The DFT time series analysis gives the wrong periods for the original data (Figure 7e: diamond at $P_{1}=0.542$ and for the residuals (Figure 7f: diamond at $P_{2}=1.059$. The DFT also gives wrong $p(t)$ trend coefficients $M_{0}=0.670$, $M_{1}=-0.025$ and $M_{2}=0.406$ (Figure 7g: dashed black line). The DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line) deviates from the data $y_{i}$ (black dots). This deviation is largest at the beginning and the end of the simulated time series. In our Figure 7h, the black dots denote the detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$, the continuous black line denotes the detrended DFT model $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ and the continuous thick red line denotes pure sine signal $s_{\mathrm{y,DFT}}(t)$ detected from the original data. The DFT model $g_{\mathrm{DFT}}(t)$ gives very low amplitude for the pure sine signal $s_{\mathrm{\epsilon,DFT}}(t)$ detected from the residuals (continuous thinner blue line). The correct simulated peak to peak amplitude for this second signal is much larger, $A_{2}=2.$ The DFT model residuals (blue dots) are offset to -1.5 level and show strong trends.
@@
The DCM analysis succeeds for Model 7 simulated time series. The DFT analysis fails, as predicted by Equations 35-38.
3.8 Summary of simulations
For all seven simulated time series, all DCM analyses succeed and all DFT analyses fail (Sections 3.1 - 3.7).
[…]
The analyses of seven simulated time series expose the DFT weaknesses. Every analysis fails! The AL6-AL10 are not the only causes for these failures. In general, the solutions for one-dimensional time series analysis models are ill-posed, if the data contain an unknown trend and/or more than one signal. For example, the detrending of simulated data misleads DFT even when the correct $K_{3}$ value is known (Sections 3.1-3.7). This correct $K_{3}$ value can be unknown for real data. The combination of detrending and iterative pre-whitening can fail at any stage. Even if the detrending succeeds, the leakage can cause the detection of wrong first signal frequency. This corrupts the model residuals and the whole analysis. For any one-dimensional time series analysis method, like the DFT, the search for many signals superimposed on an unknown trend can fail when the combination of detrending and pre-whitening is applied.
Our DCM outperforms the esteemed DTF. The DCM detects the correct signal(s) and trend, but DFT does not. For the sunspot data, the DCM detects signals superimposed on a constant trend (Jetsu 2025). The simulations presented here support the conclusion that those solar signals and the trend are correct. This would explain why the DCM is the first time series analysis method that detects Jupiter’s exact $11.^{\mathrm{y}}86$ period in the sunspot record. The one-dimensional time series analysis methods, like DFF, have failed to find exactly correct periods.
@@ 3.8 Summary of simulations
Table 8: Fisher tests. DCM analysis of Model 1 simulated time series ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.33footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
@@ 3.9 Best model
The $K_{1}$, $K_{2}$ and $K_{3}$ orders of the best model are not necessarily known when some time series analysis method is applied to the real time series. We know a priori the best DCM and DFT model orders for the simulated time series of Models 1-7 (Sections 3.1-3.7). It could be argued that our DCM analysis succeeds only for this reason.
@@
Larger $F_{R}$ or $F_{\chi}$ values have smaller $Q_{F}$ critical levels. Hence, the probability for the $H_{0}$ hypothesis rejection increases when $F_{R}$ or $F_{\chi}$ increases. If the number of complex model free parameters $\eta_{2}$ increases, the $R_{2}$ or $\chi^{2}_{2}$ values decrease. This increases the $F_{R}$ or $F_{\chi}$ values because the terms $(R_{1}/R_{2}-1)$ or $(\chi^{2}_{1}/\chi^{2}_{2}-1)$ increase (Equations 46 and 47: first terms). At the same time, the $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty term decreases (Equations 46 and 47: second terms), and this decreases the $F_{R}$ or $F_{\chi}$ values. This penalty term prevents the use of too high $\eta_{2}$ values (too complex models).
@@
We conclude that the best model for the real time series can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
3.10 Significance estimates
Jetsu [23] or Jetsu (2025) gave no signal significance estimate for the first detected period. Here, we use the Fisher-test for this purpose. The one signal model is the complex model. The logical simple model alternative is $g(t)=m\equiv$ white noise having standard deviation $s$. However, white noise is not the only possible alternative simple model. The other possible nested no signal DCM polynomial models are
$g(t)=p(t,K_{3})$ (49)
where $h(t)=0$, $K_{3}=0,1,2..$. The $K_{3}=0$ polynomial $g(t)=M_{0}$ represents white noise. The Fisher-test gives the critical level $Q_{\mathrm{F}}$ for rejecting the $H_{0}$ hypothesis when this hypothesis is in fact true (Equation 48). Therefore, this $Q_{\mathrm{F}}$ value represents the probability of false signal detection. In cases $Q_{\mathrm{F}}<\!10^{-16}$, the signal detection is absolutely certain.
We use the Model 1 simulated time series combination $n=100$ and ${\mathrm{SN}}=100$ to demonstrate the significance estimation for the first detected period. The correct model for the first detected signal is $g_{1,1,0}(t)$. The Fisher-test is used to compare this correct $g_{1,1,0}(t)$ model to different polynomial models $g(t)=p(t,K_{3})$ having $K_{3}=0,1,..,7$ (Table 9: Equation 49). The $\mathcal{M}$=1, 2 and 3 models have less free parameters than the correct $g_{1,1,0}(t)$ model, but the $\chi^{2}$ values of these three polynomial models are so large that the Fisher-test is merely a formality. The $\mathcal{M}$=4 model has the same number of free parameters as the correct $g_{1,1,0}(t)$ model, but the comparison of $\chi^{2}$ values reveals that this third order $p(t)$ polynomial is not the correct model for the time series. The next model $\mathcal{M}$=5 has more free parameters than the correct $g_{1,1,0}(t)$ model. This fourth order $p(t)$ polynomial $\mathcal{M}$=5 model must be rejected because it has a larger $\chi^{2}$ value than the correct $g_{1,1,0}(t)$ model. The critical levels $Q_{\mathrm{F}}\gg\gamma=0.001$ for the remaining $\mathcal{M}$=6, 7 and 8 polynomial models are so large that these models must also be rejected. The results in Table 9 confirm that the one signal model is better than any polynomial model in Equation 49. Hence, the analysed Model 1 simulated time series must contain at least the $h_{1}(t)$ pure sine signal. For the constant, linear or parabolic $p(t)$ trend alternatives, the significance for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
After the detection of the first strongest signal, the Fisher-test critical level $Q_{\mathrm{F}}$ values increase for the next detected weaker signals. In other words, the signal significances decrease. Typical examples can be found in Jetsu 2025 (Tables S5-S16). No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected. For the Model 1 simulated $n=100$ and ${\mathrm{SN}}=100$ combination, Table 8 shows that this time series contains only one signal.
For this Model 1 time series, there is no need to discuss the DFT significance estimates [21, their Equation 22] because this method fails to detect the correct period.
We conclude that the Fisher-test identifies the correct DCM model (Section 3.9), as well as gives the signal significance $Q_{\mathrm{F}}$ estimates (Section 3.10).
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated time series (electronic data file Model1n100SN100). (1) Polynomial model (Equation 49). (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8. 44footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.